OpenStudy (anonymous):

Algebra 2 Help!!! Simplify 1-cos^2 Θ/ cos^2 Θ A. sin^2 Θ B. cos^2 Θ C. tan^2 Θ D. 1

OpenStudy (solomonzelman):

RULE: cos^2 Θ+sin^2 Θ=1 And therefore, 1-cos^2 Θ=sin^2 Θ So, 1-cos^2 Θ/ cos^2 Θ = sin^2 Θ/ cos^2 Θ=tan^2 Θ

OpenStudy (anonymous):

Thank you so much! @SolomonZelman

OpenStudy (solomonzelman):

Anytime